OpenStudy (anonymous):

Find the solution of x2 – 14x + 3 = –10 Select one: a. {–1, 13} b. {-1, –13} c. {1, 13} d. {–1, –13}

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Make sure that all values are to the left side of the equal sign, so that you get a quadratic equation = 0

OpenStudy (anonymous):

i think its a

OpenStudy (anonymous):

x²-14x+13=0 -> D = 14²-4.13= 144 = 12² x1 = (14+12)/2 = 13 x2 = (14-12)/2 = -1 a is correct